# Critic-Guided Reinforcement Unlearning in Text-to-Image Diffusion

### v2

@@ Abstract
Machine unlearning in text-to-image diffusion models aims to remove targeted concepts while preserving overall utility. Prior diffusion unlearning methods typically rely on supervised weight edits or global penalties; reinforcement-learning (RL) approaches, while flexible, often optimize sparse end-of-trajectory rewards, yielding high-variance updates and weak credit assignment. We present a general RL framework for diffusion unlearning that treats denoising as a sequential decision process and introduces a timestep-aware critic with noisy-step rewards. Concretely, we train a CLIP-based reward predictor on noisy latents and use its per-step signal to compute advantage estimates for policy-gradient updates of the reverse diffusion kernel. Our algorithm is simple to implement, supports off-policy reuse, and plugs into standard text-to-image backbones. Across multiple concepts, the method achieves better or comparable forgetting to strong baselines while maintaining image quality and benign prompt fidelity; ablations show that (i) per-step critics and (ii) noisy-conditioned rewards are key to stability and effectiveness. We release code and evaluation scripts to facilitate reproducibility and future research on RL-based diffusion unlearning.
@@ 1 Introduction
We propose Critic-Guided Reinforcement Unlearning (CGRU), an RL framework that treats the reverse diffusion kernel as the policy and augments policy-gradient updates with a per-timestep critic. Concretely, we train a reward predictor on noisy latents to deliver timestep-aware signals  Radford et al. (2021); a learned value function over (prompt, latent, timestep) yields advantage estimates that stabilize optimization and sharpen credit assignment. CGRU supports off-policy reuse via importance weighting for sample efficiency  Kakade and Langford (2002) and integrates with standard text-to-image backbones without architectural changes. Conditioning rewards and values on intermediate latents mitigates two common failure modes in unlearning: (i) over-suppression from global penalties and (ii) instability from sparse terminal rewards.
@@ 2 Related Works
Standard policy-gradient tools motivate baselines/critics to reduce variance Greensmith et al. (2004). However, most DM alignment methods use end-of-trajectory rewards, which are sparse and high-variance, highlighting the need for timestep-aware signals.
@@
Machine unlearning aims to remove specific data or concepts from a trained model for privacy, safety, or compliance. Classic foundations include data deletion and SISA training in discriminative settings Ginart et al. (2019); Bourtoule et al. (2020). For diffusion models, concept erasure techniques target styles, identities, or unsafe content while preserving benign capabilities. Representative approaches include ESD Gandikota et al. (2023), lightweight erasers (Receler) Huang et al. (2024), closed-form editing (RECE) Gong et al. (2024), and localized/gated erasure Lee et al. (2025).
Robust forgetting remains challenging: erased concepts can often be re-elicited via adversarial prompts Tsai et al. (2024); Pham et al. (2023); Beerens et al. (2025). Evolving surveys and benchmarks continue to clarify objectives, taxonomies, and metrics in generative unlearning Feng et al. (2025).
@@ 3.1 Diffusion Models
Diffusion models learn to reverse a noise corruption process Ho et al. (2020). Given clean data $x_{0}$, the forward process adds Gaussian noise: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ where $\{\beta_{t}\}_{t=1}^{T}$ is a noise schedule. The reverse process learns to denoise by predicting noise: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}I)$ where $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ and $\epsilon_{\theta}$ is a neural network. The training objective minimizes: $\mathcal{L}=\mathbb{E}_{t,x_{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}]$.
@@ 3.2 Reinforcement Learning and Markov Decision Processes
A Markov Decision Process (MDP) is defined as $(S,A,P,R,\gamma)$ where $S$ is the state space, $A$ is the action space, $P(s^{\prime}|s,a)$ is the transition probability, $R(s,a)$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. The agent follows a policy $\pi(a|s)$ to maximize: $J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ where $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots)$ is a trajectory.
@@ 3.3 Connecting Diffusion and Reinforcement Learning
We establish a formal connection between diffusion sampling and reinforcement learning by mapping the denoising process to a Markov Decision Process. Consider a diffusion model with reverse process $p_{\theta}(x_{t-1}|x_{t})$ and a reward function $r(x_{0})$ defined on the final samples.
The MDP is defined as follows:
$s_{t}$ $=(x_{t},t)\quad\text{(state: noisy image and timestep)}$ (1)
$a_{t}$ $=x_{t-1}\quad\text{(action: denoised image)}$ (2)
$\pi_{\theta}(a_{t}|s_{t})$ $=p_{\theta}(x_{t-1}|x_{t})\quad\text{(policy: reverse process)}$ (3)
$P(s_{t+1}|s_{t},a_{t})$ $=\delta_{x_{t-1}}\otimes\delta_{t-1}\quad\text{(deterministic transition)}$ (4)
$R(s_{t},a_{t})$ $=\begin{cases}r(x_{0})&\text{if }t=0\\ 0&\text{otherwise}\end{cases}$ (5)
where $\delta_{y}$ denotes the Dirac delta distribution. The initial state distribution is $\rho_{0}(s_{0})=p(x_{T})\otimes\delta_{T}$ with $x_{T}\sim\mathcal{N}(0,I)$.
This formulation enables optimizing the diffusion model for arbitrary reward functions by maximizing:
$J(\theta)=\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}[r(x_{0})]$ (6)
The key advantage is that the policy $\pi_{\theta}(a_{t}|s_{t})=p_{\theta}(x_{t-1}|x_{t})$ remains a simple Gaussian distribution, enabling exact computation of log-likelihoods and their gradients. This is crucial for applying policy gradient methods without approximation errors.
@@ 3.4 Policy Gradient Methods for Diffusion
For the diffusion MDP, the score function estimator becomes:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim p_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t})\cdot r(x_{0})\right]$ (7)
The estimator alternates between collecting denoising trajectories and updating parameters via gradient descent. However, this approach suffers from high variance due to the sparse reward structure Mohamed et al. (2020).
However, the score function estimator requires on-policy data, limiting optimization to one step per round of data collection. To enable multiple optimization steps, we can use off-policy data through importance sampling Kakade and Langford (2002):
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim p_{\theta_{\text{old}}}}\bigg[\sum_{t=1}^{T}\frac{p_{\theta}(x_{t-1}|x_{t})}{p_{\theta_{\text{old}}}(x_{t-1}|x_{t})}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t})\cdot r(x_{0})\bigg]$ (8)
This estimator uses trajectories generated by previous parameters $\theta_{\text{old}}$ but corrects for the distribution shift through importance weights.
3.5 Actor-Critic Methods
Actor-critic methods combine an actor (policy) and critic (value function) to reduce variance in policy gradient estimation (and REINFORCE-style estimators specifically). The critic provides a learned baseline, and the advantage function A(s,a)=Q(s,a)V(s)A(s,a)=Q(s,a)-V(s) measures how much better an action is compared to the average, providing more informative signals than raw rewards.
It is worth noting that the definition of the advantage function can vary across different implementations of actor-critic methods, but the core principle remains consistent: to provide a measure of how much better or worse an action is compared to the average, thereby reducing variance in policy gradient estimation.
@@ 4.1 Problem Formulation
Following the MDP formulation from Section 3.3, we treat the reverse diffusion kernel as a policy $\pi_{\theta}(a_{t}|s_{t})=p_{\theta}(x_{t-1}|x_{t},c)$ where states are $s_{t}=(c,t,x_{t})$ and actions are $a_{t}=x_{t-1}$. The objective is to maximize:
$J(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(x_{0},c)]$ (9)
@@ 4.2 Critic-Guided Advantage Estimation
The key innovation of CGRU is the introduction of a per-timestep critic $V_{\phi}(x_{t},c,t)$ that estimates the expected terminal reward from any intermediate state:
$V_{\phi}(x_{t},c,t)\approx\mathbb{E}[r(x_{0},c)|x_{t},c,t]$ (10)
This critic enables computation of the advantage function at each timestep:
$A(s_{t},a_{t})=r(x_{0},c)-V_{\phi}(x_{t},c,t)$ (11)
The advantage function measures how much better (or worse) the current state is compared to the expected outcome, providing more informative signals than the raw terminal reward.
@@ 4.3 Advantage-Weighted Policy Updates with Importance Sampling
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau\sim p_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot A(s_{t},a_{t})\right]$ (12)
To enable off-policy learning and sample efficiency, we extend this to importance sampling (Sec 3.4). When using trajectories generated by a previous policy $p_{\theta_{\text{old}}}$, the gradient becomes:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau\sim p_{\theta_{\text{old}}}}\left[\sum_{t=1}^{T}\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\theta_{\text{old}}}(x_{t-1}|x_{t},c)}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot A(s_{t},a_{t})\right]$ (13)
This formulation maintains the same optimization objective as standard policy gradient methods but with much lower variance due to the learned baseline and enables efficient reuse of previously collected trajectories. The complete proof is provided in Appendix A.
@@ 4.4 Practical Implementation
The CGRU framework consists of two main training phases: critic training and policy optimization. We describe the practical implementation details below.
4.4.1 Reward Functions and Datasets
The choice of reward function and training datasets is crucial for effective unlearning. The reward function $r(x_{0},c)$ should penalize the presence of unwanted concepts while preserving overall image quality and prompt adherence. Common approaches include CLIP-based similarity measures to detect concept presence, classifier outputs for specific unwanted content, or other domain-specific metrics depending on the target concept.
The training datasets typically consist of two components: a forgetting dataset containing prompts that directly elicit the concept to be removed, and a retention dataset with diverse prompts that should continue to work normally. The contents, relative composition and size of these components can be adjusted based on the specific unlearning requirements.
4.4.2 Critic Training
The critic $V_{\phi}(x_{t},c,t)$ is trained to predict expected terminal reward from intermediate states through supervised learning on generated trajectories. Training involves: (1) generating denoising trajectories $(x_{T},x_{T-1},\ldots,x_{0})$ for each prompt, (2) computing terminal rewards $r(x_{0},c)$, (3) shuffling timesteps to break temporal dependencies, and (4) minimizing:
$\mathcal{L}_{\text{critic}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\left\|V_{\phi}\!\left(x_{t}^{(i)},c^{(i)},t\right)-r\!\left(x_{0}^{(i)},c^{(i)}\right)\right\|^{2}$ (14)
The critic uses sinusoidal timestep embeddings and FiLM layers Perez et al. (2017) for temporal awareness, enabling meaningful value estimates across denoising timesteps. Details are in Appendix C.
4.4.3 Policy Optimization
Policy optimization uses advantage-weighted gradients: (1) compute advantages $A_{t}=r(x_{0},c)-V_{\phi}(x_{t},c,t)$ at every timestep, (2) shuffle timesteps to break temporal dependencies, (3) apply importance weighting for off-policy data, and (4) update the policy with gradient clipping, following DDPO Black et al. (2024).
Complete algorithms are in Appendix B.
@@ 5.1 Experimental Setup
All experiments were conducted using Stable Diffusion 1.5 as the base diffusion model, with LoRA (Low-Rank Adaptation) applied for memory efficiency during fine-tuning. Training was performed using gradient accumulation with 2 updates per epoch, requiring approximately 35 GB of VRAM on a single H100 GPU. Each training run completed in approximately 2 hours. Detailed hyperparameters and training configurations are provided in Appendix D.
@@ 5.2 Comparison with DDPO
We compare CGRU against DDPO Black et al. (2024) using the Aesthetic Score reward function. Although DDPO is not a specialized unlearning method, it serves as an ablation baseline to validate our theoretical claims regarding the superiority of dense, timestep-aware signals over sparse terminal rewards. Both methods use identical configurations, with the only difference being CGRU’s per-timestep critic versus DDPO’s terminal rewards.
@@
The improved performance of CGRU validates our theoretical analysis, suggesting that the per-timestep critic provides more effective optimization signals compared to sparse, end-of-trajectory rewards. Additionally, we observe that CGRU exhibits lower gradient variance during training (see Appendix E.2), confirming the stability benefits of our approach. This result demonstrates the practical viability of CGRU for RL-based diffusion model optimization.

### v3

@@ Abstract
Machine unlearning in text-to-image diffusion models aims to remove targeted concepts while preserving overall utility. Prior diffusion unlearning methods typically rely on supervised weight edits or global penalties; reinforcement-learning (RL) approaches, while flexible, often optimize sparse end-of-trajectory rewards, yielding high-variance updates and weak credit assignment. We present a general RL framework for diffusion unlearning that treats denoising as a sequential decision process and introduces a timestep-aware critic that predicts expected terminal reward from noisy intermediate states. Concretely, we train a CLIP-based predictor on noisy intermediate states and use it to estimate per-timestep values to compute advantage estimates for policy-gradient updates of the reverse diffusion kernel. Our algorithm is simple to implement, supports off-policy reuse, and plugs into standard text-to-image backbones. Across multiple concepts, the method achieves better or comparable forgetting to strong baselines while maintaining image quality and benign prompt fidelity; ablations show that (i) per-step critics and (ii) noisy-conditioned value estimates are key to stability and effectiveness. We release code and evaluation scripts to facilitate reproducibility and future research on RL-based diffusion unlearning.
@@ 1 Introduction
We propose Critic-Guided Reinforcement Unlearning (CGRU), an RL framework that treats the reverse diffusion kernel as the policy and augments policy-gradient updates with a per-timestep critic. Concretely, we train a reward predictor on noisy latents to deliver timestep-aware signals; a learned value function over (prompt, latent, timestep) yields advantage estimates that stabilize optimization and sharpen credit assignment. CGRU supports off-policy reuse via importance weighting for sample efficiency  Kakade and Langford (2002) and integrates with standard text-to-image backbones without architectural changes. Conditioning rewards and values on intermediate latents mitigates two common failure modes in unlearning: (i) over-suppression from global penalties and (ii) instability from sparse terminal rewards.
@@ 2 Related Works
Standard policy-gradient tools motivate baselines/critics to reduce variance Greensmith et al. (2004). However, most DM alignment methods use end-of-trajectory rewards, which are sparse and high-variance, highlighting the need for timestep-aware signals. Recent work continues to improve sample efficiency and stability for RL-based diffusion fine-tuning Gupta et al. (2025).
@@
Machine unlearning aims to remove specific data or concepts from a trained model for privacy, safety, or compliance. Classic foundations include data deletion and SISA training in discriminative settings Ginart et al. (2019); Bourtoule et al. (2020). For diffusion models, concept erasure techniques target styles, identities, or unsafe content while preserving benign capabilities. Representative approaches include ESD Gandikota et al. (2023), lightweight erasers (Receler) Huang et al. (2024), closed-form editing (RECE) Gong et al. (2024), and localized/gated erasure Lee et al. (2025a).
Robust forgetting remains challenging: erased concepts can often be re-elicited via adversarial prompts Tsai et al. (2024); Pham et al. (2023); Beerens et al. (2025). Evolving surveys and benchmarks continue to clarify objectives, taxonomies, and metrics in generative unlearning Feng et al. (2025); notably, the Holistic Unlearning Benchmark (HUB) Moon et al. (2025) proposes multi-faceted evaluation across faithfulness, robustness, and alignment dimensions.
@@ 3.1 Diffusion Models
Diffusion models learn to reverse a noise corruption process Ho et al. (2020). Given clean data $x_{0}$, the forward process adds Gaussian noise: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ where $\{\beta_{t}\}_{t=1}^{T}$ is a noise schedule. The reverse process learns to denoise by predicting noise: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}I)$ where $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ and $\epsilon_{\theta}$ is a neural network. The training objective minimizes: $\mathcal{L}=\mathbb{E}_{t,x_{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}]$. Here $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$.
@@ 3.2 Reinforcement Learning and Markov Decision Processes
A Markov Decision Process (MDP) is defined as $(S,A,P,R,\gamma)$ where $S$ is the state space, $A$ is the action space, $P(s^{\prime}|s,a)$ is the transition probability, $R(s,a)$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. The agent follows a policy $\pi(a|s)$ to maximize: $J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ where $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots)$ is a trajectory. In our diffusion-as-RL formulation, we use a finite-horizon episodic MDP of length $T$ with a terminal reward at the end of the denoising trajectory.
@@ 3.3 Connecting Diffusion and Reinforcement Learning
We cast reverse diffusion as a sequential decision process: the noisy latent and timestep form the state, the sampled previous latent $x_{t-1}$ is the action, and the reverse kernel acts as the policy. Rewards are assessed on the final sample. This view lets us apply standard policy-gradient tools to diffusion without architectural changes. Full formal definitions and derivations are provided in Appendix A.
@@ 3.4 Policy Gradient Methods for Diffusion
For the diffusion MDP, by Mohamed et al. (2020), the score function estimator becomes:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim p_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t})\cdot r(x_{0})\right]$
Additionally, we adopt importance sampling Kakade and Langford (2002), reusing off-policy trajectories, to enable multiple optimization steps.
@@ 4.1 Problem Formulation
Following the MDP formulation in Appendix A, we treat the reverse diffusion kernel as a policy $\pi_{\theta}(a_{t}|s_{t})=p_{\theta}(x_{t-1}|x_{t},c)$ where states are $s_{t}=(c,t,x_{t})$ and actions are $a_{t}=x_{t-1}$. The objective is to maximize:
$J(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(x_{0},c)]$
@@ 4.2 Critic-Guided Advantage Estimation
The key innovation of CGRU is the introduction of a per-timestep critic $V_{\phi}(x_{t},c,t)$ that estimates the expected terminal reward from any intermediate state. This critic enables computation of the advantage function $A(s_{t})=r(x_{0},c)-V_{\phi}(x_{t},c,t)$ at each timestep. In actor–critic terms, $V_{\phi}$ serves as a learned baseline that reduces variance of REINFORCE-style gradients by centering terminal rewards at each state; the resulting advantage weights provide more informative signals than raw rewards  Kakade and Langford (2002).
@@ 4.3 Advantage-Weighted Policy Updates with Importance Sampling
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau\sim p_{\theta}}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot A(s_{t})\right]$
In practice, we reuse trajectories within the same outer iteration: after collecting a batch of $N$ trajectories with policy $\theta_{\mathrm{old}}$, we perform $K$ gradient updates on this batch. We do not maintain a long-lived replay buffer across many past policies; “off-policy” refers to updating $\theta$ using data sampled from $\theta_{\mathrm{old}}$ via importance weighting Kakade and Langford (2002). We monitor importance ratios and stabilize training with gradient clipping; in our settings, ratios remain well-behaved (see Appendix B).
We additionally clip importance ratios to $\rho_{\min},\rho_{\max}]$ (PPO-style) Schulman et al. (2017) for stability; this does not change our conclusions (Appendix B).
@@ 4.4 Practical Implementation
The CGRU framework consists of two main components: critic learning (warm-start + online updates) and policy optimization. We describe the practical implementation details below.
Reward Functions and Datasets.
The choice of reward function and training datasets is crucial for effective unlearning. The reward function $r(x_{0},c)$ should penalize the presence of unwanted concepts while preserving overall image quality and prompt adherence. Common approaches include CLIP-based similarity measures to detect concept presence (applied in image space after decoding latents), classifier outputs for specific unwanted content, or other domain-specific metrics depending on the target concept. Importantly, the critic is a standalone module that can be warm-started independently of the diffusion model’s fine-tuning process; in our implementation, we continue training it online during policy optimization to reduce distribution shift as the diffusion model evolves.
The training dataset consists of a forgetting dataset containing prompts that directly elicit the concept to be removed. We train only on this forgetting dataset in our experiments.
Critic Training.
The critic $V_{\phi}(x_{t},c,t)$ is trained to predict expected terminal reward from intermediate states through supervised learning on generated trajectories, and it keeps training during diffusion model fine-tuning. Training involves: (1) generating denoising trajectories $(x_{T},x_{T-1},\ldots,x_{0})$ for each prompt, (2) computing terminal rewards $r(x_{0},c)$, (3) shuffling timesteps to break temporal dependencies, and (4) minimizing the MSE loss:
$\mathcal{L}_{\text{critic}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\left\|V_{\phi}\!\left(x_{t}^{(i)},c^{(i)},t\right)-r\!\left(x_{0}^{(i)},c^{(i)}\right)\right\|^{2}$
In our implementation, $V_{\phi}$ is instantiated with a CLIP-based vision classifier: each intermediate state $x_{t}$ (a noisy latent) is decoded to RGB space using the frozen VAE decoder, and the resulting image is fed to the CLIP vision encoder with a lightweight classification head.
The critic uses sinusoidal timestep embeddings and FiLM layers Perez et al. (2017) for temporal awareness, enabling meaningful value estimates across denoising timesteps. Details are in Appendix E.
Policy Optimization.
Policy optimization uses advantage-weighted gradients: (1) compute advantages $A_{t}=r(x_{0},c)-V_{\phi}(x_{t},c,t)$ at every timestep, (2) shuffle timesteps to break temporal dependencies (as in critic training), (3) apply importance weighting for off-policy data, and (4) update the policy with stability controls. To ensure stable off-policy updates, we employ gradient clipping, following DDPO Black et al. (2024); these mechanisms constrain the policy update magnitude and prevent excessively large steps that could destabilize training.
All details on objective derivations and proofs are available in Appendix A and complete algorithms are in Appendix D.
@@ 5.1 Experimental Setup
All experiments were conducted using Stable Diffusion 1.5 as the base diffusion model, with LoRA (Low-Rank Adaptation) applied for memory efficiency during fine-tuning. Training was performed using gradient accumulation with 2 updates per epoch, requiring approximately 35 GB of VRAM on a single H100 GPU. Each training run took approximately 2–5 hours (100–250 epochs), depending on the target class and its convergence speed. Detailed hyperparameters and training configurations are provided in the Appendix C. We aggregate results over 3 random seeds.
@@ 5.2 Comparison with DDPO
We compare CGRU against DDPO Black et al. (2024) using the LAION Aesthetic Score Schuhmann (2022) reward function - aesthetics predictor, built as a linear layer applied to CLIP embeddings. Although DDPO is not a specialized unlearning method, it serves as an ablation baseline to validate our theoretical claims regarding the superiority of dense, timestep-aware signals over sparse terminal rewards. Both methods share the same base model, reward, dataset, etc. The key difference is the learning algorithm itself: CGRU augments the update with a per-timestep critic (enabling dense advantages and trajectory reuse), whereas DDPO follows the standard on-policy baseline with terminal rewards.
@@
The improved performance of CGRU validates our theoretical analysis, suggesting that the per-timestep critic provides more effective optimization signals compared to sparse, end-of-trajectory rewards. Additionally, we observe that CGRU exhibits a lower gradient variance during training (see Appendix B.4), which confirms the stability benefits of our approach. This result demonstrates the practical viability of CGRU for the optimization of RL-based diffusion models.